OpenStudy (anonymous):

Find f multiplied by g, and f divided by g. State any domain restrictions. f(x) = -x^2; g(x) = 3x. Can someone please explain this to me? I don't understand a thing!

OpenStudy (anonymous):

f*g no restrciction in the domain, f*g=-3*x^3

OpenStudy (anonymous):

f/g=-x²/3x, but can we divide a number by zero? so g(x), must be different form zero this domain will be\[x \in R |x \neq 0\]

OpenStudy (anonymous):

f/g= -x/3

OpenStudy (anonymous):

@__mia_rachel__ get it?

OpenStudy (anonymous):

Not at all.

OpenStudy (anonymous):

HELP

OpenStudy (anonymous):

what you dont get?

OpenStudy (anonymous):

I don't understand ANYTHING. my teacher is horrible!

OpenStudy (anonymous):

I don't even know what a function is -_-

OpenStudy (anonymous):

for example f(x)= x² this function associate every number another number in this case the law of formation is this you pick a number and square it

OpenStudy (anonymous):

the domain is all possible x that i can calculate that function

OpenStudy (anonymous):

Whhaaatttt???

OpenStudy (anonymous):